# Isometries of 3-dimensional semi-Riemannian Lie groups

### v1

@@ Abstract.
Let $G$ be a connected, simply connected three-dimensional Lie group (unimodular or non-unimodular) equipped with a left-invariant (Riemannian or Lorentzian) metric $g$. Using Lie theoretical methods, we present a unified approach to obtain all pairs $(G,g)$ whose full isometry group $\mathrm{Isom}(G,g)$ has dimension greater or equal to four. As a consequence, we determine, for every pair $(G,g)$ up to automorphism and scaling, the dimension of $\mathrm{Isom}(G,g)$, which can be three, four or six.
@@ 1. Introduction
The isometry group of a semi-Riemannanian manifold has the structure of a Lie group in a natural way ([O’N83, Ch. 9, Th. 32]). For compact Riemannian manifolds, it is well-known that the isometry group is compact [MS39]. For semi-Riemannian manifolds this is not always the case, however, proved to be true for Lorentzian simply connected, compact (real analytic) manifolds, see [D’A88] and examples therein. In a recent publication [ZCL21], it was shown that the isometry group of a left-invariant semi-Riemannian metric of a connected, simply connected, compact simple Lie group is compact.
@@
In particular, for a 3-dimensional metric Lie group, the dimension of the full isometry group is between 3 and 6. It is known that this dimension cannot be 5 (cf. for instance [ABZ23]), so the only possibilities are 3, 4, or 6.
Many articles have been devoted to the study of isometries of left-invariant metrics on Lie groups of dimension 3. We have the seminal article of Milnor, [Mil76], where the list of the 3-dimensional unimodular Lie algebras was derived, and curvature properties, such as constant sectional curvature, of Riemannian metrics where studied. In [Shi97], a characterization, which turns out to be independent of curvature properties, of the isometry group of left-invariant Riemannian metrics on each of the six unimodular connected simply connected 3-dimensional Lie groups is given. The same question is considered in [HL12], where a different approach is used relying on normal forms of metrics and principal Ricci curvatures. Only recently, as far as the authors know, there were developments in the Lorentzian or in the non-unimodular Lie group cases. More precisely, a similar approach to that of [HL12] was employed in [BC22a] to determine the isometry groups of all 3-dimensional, connected, simply connected and unimodular Lie groups endowed with a left-invariant Lorentzian metric. On the other hand, in [CR22], the full isometry group of any Riemannian left-invariant metric on a simply connected, non-unimodular Lie group of dimension 3 was computed.
@@ 1.1. Bianchi classification
For each connected and simply connected Lie group $G$, the left-invariant metrics will be given at the Lie algebra level $\mathfrak{g}$ using the basis of 1-forms dual to the corresponding preferred basis of Table 1. Furthermore, the metrics will be given in normal form, that is, we will present a representative of a given metric in their orbit under the action of ×$\mathbb{R}^{*}\times\mathrm{Aut}(\mathfrak{g})$. Observe that by identifying the metrics whose isometry Lie algebra is of dimension $6$, Table 2, and of dimension $4$, Table 3, we are automatically identifying the metrics whose isometry Lie algebra is of dimension 3, as they are the remaining ones. The list of normal forms of metrics for each Lie algebra can be found in Appendix A.
Lie algebra Non-vaninishing brackets Automorphism group
@@
$\mathfrak{so}(3)$ $[e_{1},e_{2}]=e_{3},[e_{2},e_{3}]=e_{1},[e_{3},e_{1}]=e_{2}$ $\mathrm{SO}(3)$
$\mathfrak{sl}(2,\mathbb{R})$ $[e_{1},e_{2}]=e_{2},[e_{3},e_{1}]=e_{3},[e_{2},e_{3}]=e_{1}$ $\mathrm{SO}(2,1)$
$\mathfrak{heis}$ $[e_{1},e_{2}]=e_{3}$$\left\{\left(\begin{smallmatrix}a&b&0\\ c&d&0\\ e&f&ad-bc\end{smallmatrix}\right):{\tiny ad-bc\neq 0}\right\}$
$\mathfrak{euc}(2)$ $[e_{1},e_{2}]=e_{3},[e_{3},e_{1}]=e_{2}$$\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&-c\\ d&c&b\end{smallmatrix}\right),\left(\begin{smallmatrix}-1&0&0\\ a&-b&c\\ d&c&b\end{smallmatrix}\right):{\tiny b^{2}+c^{2}\neq 0}\right\}$
$\mathfrak{sol}$ $[e_{1},e_{2}]=e_{2},[e_{3},e_{1}]=e_{3}$$\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&0\\ c&0&d\end{smallmatrix}\right),\left(\begin{smallmatrix}-1&0&0\\ a&0&b\\ c&d&0\end{smallmatrix}\right):{\tiny bd\neq 0}\right\}$
$\mathfrak{aff}(\mathbb{R})\oplus\ \mathbb{R}$ $[e_{1},e_{2}]=e_{2}$$\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&0\\ c&0&d\end{smallmatrix}\right):{\tiny bd\neq 0}\right\}$
$\mathfrak{h}(1)$ $[e_{1},e_{2}]=e_{2},[e_{1},e_{3}]=e_{3}$$\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&c\\ d&e&f\end{smallmatrix}\right):{\tiny bf-ce\neq 0}\right\}$
$\mathfrak{psh}$ $[e_{1},e_{2}]=e_{2},[e_{1},e_{3}]=e_{2}+e_{3}$$\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&c\\ d&0&b\end{smallmatrix}\right):{\tiny b\neq 0}\right\}$
[…]
$\mathfrak{e}(\mu),\,\mu>0$ $[e_{1},e_{2}]=\mu e_{2}+e_{3},[e_{1},e_{3}]=\mu e_{3}-e_{2}$$\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&-c\\ d&c&b\end{smallmatrix}\right):{\tiny b^{2}+c^{2}\neq 0}\right\}$
@@ 1.2. Strategy and methods
The linear algebra approach which we described in the previous subsection has some important subtleties. If $I$ is the isotropy subgroup at $e_{G}$ and we identify $G$ with the left translations $L(G)$, then the full isometry group $L$ is given by $L=I.G$, but there is no straighforward way of determining the algebraic structure of $L$ from the ones of $I$ and $G$. In particular, $I.G$ is not necessarily a semidirect product. Furthermore, $G$ is normal in $L$ if and only if the connected component of the identity in $I$ is contained in $\mathrm{Aut}(\mathfrak{g})$, see [Shi97] or Subsec. 2.3.
Now, suppose we start with the algebraic data $(\mathfrak{L},\mathfrak{g},\mathfrak{I})$. From the transversality assumption $\mathfrak{g}\oplus\mathfrak{I}=\mathfrak{L}$, it is only possible to know a priori that $G$ acts on $L/I$ with an open orbit. A crucial point is that the purely algebraic approach can also capture incomplete Killing vector fields which only integrate to local isometries. For instance, in Section 5, we present the example of the product of 2-dimensional de Sitter space with $\mathbb{R}$ $,(\mathrm{dS_{2}},ds^{2})\times(\mathbb{R},dt^{2})$, which satisfies all of our assumptions but it does not have, nonetheless, a transitive action of any 3-dimensional Lie group. In section 7, we prove that the action of $G$ on $L/I$ is necessarily transitive except in the following two cases: (i) $[\mathfrak{L},\mathfrak{L}]=\mathbb{R}^{2}\not\subset\mathfrak{g}$ and (ii) $\mathfrak{g}$ is solvable but $\mathfrak{L}$ is reductive. It is interesting to observe that these exceptional cases always appear for $\mathfrak{g}=\mathrm{aff}(\mathbb{R})\oplus\mathbb{R}$. In the cases where transitivity is not guaranteed, then further considerations are required, and the necessary analysis is provided also in Sec. 7.
@@ 1.4. Main results
The contents of this article can be neatly summarized in Tables 2 and 3. When we set out to determine which metric Lie groups have 4-dimensional isometry group, we will naturally encounter the space forms, since the 6-dimensional full isometry Lie algebra of a space form will have 4-dimensional subalgebras. We identify them by computing their curvature tensors and observing if they are Einstein (in dimension 3, this is equivalent to constant sectional curvature) and checking if the metric is geodesically complete. For the completeness results, we make use of the work of Broomberg and Medina [BM08] in the unimodular case and of two other articles in our program [CFZ24, CFZ25]. It is well-known that, for a fixed dimension, signature and sectional curvature, all space forms are isometric [O’N83, Ch. 8, Prop. 23]. The question of which metric Lie groups realize space forms is a natural one. The flat case has been considered by Boucetta and Lebzioui, [BL16]. In the 3-dimensional case, Table 2 presents (up to automorphism and scaling) a full answer to this question.
@@
Table 3 presents all the metric connected and simply connected Lie groups (represented by their Lie algebras) whose isometry group is exactly of dimension 4. We add the isotropy type in each case, as well as if $\mathfrak{g}$ is an ideal in $\mathfrak{L}$ or not. Also, we include the derived subalgebra of $\mathfrak{L}$, $[\mathfrak{L},\mathfrak{L}]$, so that the subsection where each metric appears can be easily identified.
$\mathfrak{g}$ Metric normal forms Isotropy type [$\mathfrak{L}$, $\mathfrak{L}$] $\mathfrak{g}$ ideal in $\mathfrak{L}$
@@
$\mathrm{aff}(\mathbb{R})\oplus\mathbb{R}$ $(e^{1})^{2}+2(e^{2}e^{3})$ hyperbolic $\mathbb{R}^{2}$ yes
@@
Some comments about Table 3 are in order. We start by observing that $\mathfrak{euc}(2)$ does not appear in this table, there are no metrics of 1-dimensional isotropy. The unique metrics on $\mathfrak{psh}$ and on each $\mathfrak{e}(\mu)$ with $\mu>0$ are of constant sectional curvature (zero on $\mathfrak{psh}$ and positive on $\mathfrak{e}(\mu)$). On $\mathfrak{h}(1)$, it is known that all Lorentzian metrics are incomplete, [VS20]. They are all Einstein and apart from the Riemannian (which corresponds to hyperbolic space $\mathbb{H}^{3})$, all have isometry group of dimension 4. It can be seen from Appendix A that all metric normal forms on $\mathfrak{h}(1)$ are present in Tables 2 and 3. A similar situation happens for the Heisenberg group, the lists contains all normal forms of metrics. We have a Lorentzian flat space form, and all the other metrics have isometry group of dimension 4. However, all metrics on $\mathfrak{heis}$ are complete, since $\mathfrak{heis}$ is 2-step nilpotent [Gue94].
@@ 2. Preliminaries
Let $(G,g)$ be a 3-dimensional metric Lie group, that is, a Lie group $G$ equipped with a left-invariant (Riemannian or Lorentzian) metric $g$. We will assume that $G$ is connected and simply connected (see [O’N83, Ch. 9, Prop. 20]). Let $L$ denote the full isometry group of $(G,g)$ and $I$ the isotropy subgroup of $L$ at the identity $e_{G}\in G$. Denote also by $\mathfrak{g}$, $\mathfrak{L}$, and $\mathfrak{I}$ the Lie algebras of $G$, $L$ and $I$, respectively. We shall be referring to $\mathfrak{L}$ as the Killing algebra of $(G,g)$ and to its derived subalgebra $[\mathfrak{L},\mathfrak{L}]$ as the derived Killing algebra.
@@ 2.1. Infinitesimal Lorentzian isometries
Let $(V,\langle-,-\rangle)$ be a linear Lorentzian space, that is, a vector space $V$ of dimension $n$ equipped with a scalar product of signature $(n-1,1)$.
@@
A linear endomorphism $A:V\longrightarrow V$ is said to be an infinitesimal Lorentzian isometry if $A$ is skew-symmetric with respect to $\langle$-,$\rangle$ - , - ⟩, i.e.
@@ 2.1.1. Hyperbolic type
$A$ has 3 real eigenvalues not all vanishing.
@@ 2.1.2. Parabolic (or nilpotent) type
Then $A$ is nilpotent. It can be readily shown, from the non-degeneracy of $\langle$-,$\rangle$ - , - ⟩ that the rank of $A$ cannot be 1. This means that A cannot be 2-nilpotent. Therefore there exists $u\in V$ such that $\{A^{2}(u),A(u),u\}$ is a basis of $V$ and in which the matrix of $A$ is written as $\left(\begin{smallmatrix}0&1&0\\ 0&0&1\\ 0&0&0\end{smallmatrix}\right)$. The exponential of $A$, $\mathrm{exp(A)}$, is unipotent and has a fixed lightlike direction $\ell$ (the span of $A^{2}(u)$) and, therefore, also fixes the orthogonal lightlike plane to $\ell$ (which contains $\ell$).

### v2

@@ Abstract.
Let $G$ be a connected, simply connected three-dimensional Lie group (unimodular or non-unimodular) equipped with a left-invariant (Riemannian or Lorentzian) metric $g$. By definition, the isometry group $\mathrm{Isom}(G,g)$ contains $G$ itself, acting by left translations. It turns out that, generically, $\mathrm{Isom}(G,g)$ is actually equal to $G$, and the natural question then becomes to classify those special metrics for which this is not the case. Using Lie-theoretical methods, we present a unified approach to obtain all pairs $(G,g)$ whose full isometry group $\mathrm{Isom}(G,g)$ has dimension greater than or equal to four. As a consequence, we determine, for every pair $(G,g)$, up to automorphism and scaling, the dimension of $\mathrm{Isom}(G,g)$, which can be three, four, or six.
Keywords: Isometries, Killing Lie algebra, left-invariant metrics, isotropy representation.
@@
Corresponding author: Ana Cristina Ferreira (anaferreira@math.uminho.pt)
@@ 1. Introduction
The isometry group of a semi-Riemannian manifold has the structure of a Lie group in a natural way ([O’N83, Ch. 9, Th. 32]). For compact Riemannian manifolds, it is well-known that the isometry group is compact [MS39]. For semi-Riemannian manifolds this is not always the case, however, proved to be true for Lorentzian simply connected, compact (real analytic) manifolds, see [D’A88] and examples therein. In a recent publication [ZCL21], it was shown that the isometry group of a left-invariant semi-Riemannian metric of a connected, simply connected, compact simple Lie group is compact.
@@
In particular, for a 3-dimensional metric Lie group, the dimension of the full isometry group is between 3 and 6. It is known that this dimension cannot be 5 (see, for instance, [ABZ23]), so the only possibilities are 3, 4, or 6.
Many articles have been devoted to the study of isometries of left-invariant metrics on Lie groups of dimension 3. In the seminal article of Milnor, [Mil76], the list of 3-dimensional unimodular Lie algebras was derived, and curvature properties, such as constant sectional curvature, of Riemannian metrics were studied. In [Shi97], a characterization, which turns out to be independent of curvature properties, of the isometry group of left-invariant Riemannian metrics on each of the six unimodular connected simply connected 3-dimensional Lie groups is given. The same question is considered in [HL12], where a different approach is used relying on normal forms of metrics and principal Ricci curvatures. Only recently, as far as the authors know, there were developments in the Lorentzian or in the non-unimodular Lie group cases. More precisely, a similar approach to that of [HL12] was employed in [BC22a] to determine the isometry groups of all 3-dimensional, connected, simply connected and unimodular Lie groups endowed with a left-invariant Lorentzian metric. On the other hand, in [CR22], the full isometry group of any Riemannian left-invariant metric on a simply connected, non-unimodular Lie group of dimension 3 was computed.
@@ 1.1. Bianchi classification
For each connected and simply connected Lie group $G$, the left-invariant metrics will be given at the Lie algebra level $\mathfrak{g}$ using the basis of 1-forms dual to the corresponding preferred basis of Table 1. Furthermore, the metrics will be given in normal form, that is, we will present a representative of a given metric in their orbit under the action of $\mathbb{R}^{*}\times\mathrm{Aut}(\mathfrak{g})$. Observe that by identifying the metrics whose isometry Lie algebra is of dimension $6$, Table 2, and of dimension $4$, Table 3, we are automatically identifying the metrics whose isometry Lie algebra is of dimension 3, as they are the remaining ones. The list of normal forms of metrics for each Lie algebra can be found in Appendix A.
Lie algebra Non-vanishing brackets Automorphism group
@@
$\mathfrak{so}(3)$ $e_{1},e_{2}]=e_{3},[e_{2},e_{3}]=e_{1},[e_{3},e_{1}]=e_{2}$ $\mathrm{SO}(3)$
$\mathfrak{sl}(2,\mathbb{R})$ $e_{1},e_{2}]=e_{2},[e_{3},e_{1}]=e_{3},[e_{2},e_{3}]=e_{1}$ $\mathrm{SO}(2,1)$
$\mathfrak{heis}$ $e_{1},e_{2}]=e_{3}$ $\left\{\left(\begin{smallmatrix}a&b&0\\ c&d&0\\ e&f&ad-bc\end{smallmatrix}\right):{\tiny ad-bc\neq 0}\right\}$
$\mathfrak{euc}(2)$ $e_{1},e_{2}]=e_{3},[e_{3},e_{1}]=e_{2}$ $\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&-c\\ d&c&b\end{smallmatrix}\right),\left(\begin{smallmatrix}-1&0&0\\ a&-b&c\\ d&c&b\end{smallmatrix}\right):{\tiny b^{2}+c^{2}\neq 0}\right\}$
$\mathfrak{sol}$ $e_{1},e_{2}]=e_{2},[e_{3},e_{1}]=e_{3}$ $\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&0\\ c&0&d\end{smallmatrix}\right),\left(\begin{smallmatrix}-1&0&0\\ a&0&b\\ c&d&0\end{smallmatrix}\right):{\tiny bd\neq 0}\right\}$
$\mathfrak{aff}(\mathbb{R})\oplus\ \mathbb{R}$ $e_{1},e_{2}]=e_{2}$ $\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&0\\ c&0&d\end{smallmatrix}\right):{\tiny bd\neq 0}\right\}$
$\mathfrak{h}(1)$ $e_{1},e_{2}]=e_{2},[e_{1},e_{3}]=e_{3}$ $\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&c\\ d&e&f\end{smallmatrix}\right):{\tiny bf-ce\neq 0}\right\}$
$\mathfrak{psh}$ $e_{1},e_{2}]=e_{2},[e_{1},e_{3}]=e_{2}+e_{3}$ $\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&c\\ d&0&b\end{smallmatrix}\right):{\tiny b\neq 0}\right\}$
$\mathfrak{h}(\lambda),\,0<|\lambda|<1$ $e_{1},e_{2}]=e_{2},[e_{1},e_{3}]=\lambda e_{3}$ $\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&0\\ c&0&d\end{smallmatrix}\right):{\tiny bd\neq 0}\right\}$
$\mathfrak{e}(\mu),\,\mu>0$ $e_{1},e_{2}]=\mu e_{2}+e_{3},[e_{1},e_{3}]=\mu e_{3}-e_{2}$ $\left\{\left(\begin{smallmatrix}1&0&0\\ a&b&-c\\ d&c&b\end{smallmatrix}\right):{\tiny b^{2}+c^{2}\neq 0}\right\}$
@@ 1.2. Strategy and methods
The linear algebra approach which we described in the previous subsection has some important subtleties. If $I$ is the isotropy subgroup at $e_{G}$ and we identify $G$ with the left translations $L(G)$, then the full isometry group $L$ is given by $L=I.G$, but there is no straightforward way of determining the algebraic structure of $L$ from the ones of $I$ and $G$. In particular, $I.G$ is not necessarily a semidirect product. Furthermore, $G$ is normal in $L$ if and only if the connected component of the identity in $I$ is contained in $\mathrm{Aut}(\mathfrak{g})$, see [Shi97] or Subsec. 2.3.
Now, suppose we start with the algebraic data $(\mathfrak{L},\mathfrak{g},\mathfrak{I})$. From the transversality assumption $\mathfrak{g}\oplus\mathfrak{I}=\mathfrak{L}$, it is only possible to know a priori that $G$ acts on $L/I$ with an open orbit. A crucial point is that the purely algebraic approach can also capture incomplete Killing vector fields which only integrate to local isometries. For instance, in Section 5, we present the example of the product of 2-dimensional de Sitter space with $\mathbb{R}$ $(\mathrm{dS_{2}},ds^{2})\times(\mathbb{R},dt^{2}),$ which satisfies all of our assumptions but it does not have, nonetheless, a transitive action of any 3-dimensional Lie group. In section 7, we prove that the action of $G$ on $L/I$ is necessarily transitive except in the following two cases: (i) $\mathfrak{L},\mathfrak{L}]=\mathbb{R}^{2}\not\subset\mathfrak{g}$ and (ii) $\mathfrak{g}$ is solvable but $\mathfrak{L}$ is reductive. It is interesting to observe that these exceptional cases always appear for $\mathfrak{g}=\mathrm{aff}(\mathbb{R})\oplus\mathbb{R}$. In the cases where transitivity is not guaranteed, then further considerations are required, and the necessary analysis is provided also in Sec. 7.
@@ 1.4. Main results
The contents of this article can be neatly summarized in Tables 2 and 3. When we set out to determine which metric Lie groups have 4-dimensional isometry group, we will naturally encounter the space forms, since the 6-dimensional full isometry Lie algebra of a space form will have 4-dimensional subalgebras. We identify them by computing their curvature tensors and observing if they are Einstein (in dimension 3, this is equivalent to constant sectional curvature) and checking if the metric is geodesically complete. For the completeness results, we make use of the work of Broomberg and Medina [BM08] in the unimodular case and of two other articles in our program [CFZ24, CF25]. It is well-known that, for a fixed dimension, signature and sectional curvature, all space forms are isometric [O’N83, Ch. 8, Prop. 23]. The question of which metric Lie groups realize space forms is a natural one. The flat case has been considered by Boucetta and Lebzioui, [BL16]. In the 3-dimensional case, Table 2 presents (up to automorphism and scaling) a full answer to this question.
@@
Table 3 presents all the metric connected and simply connected Lie groups (represented by their Lie algebras) whose isometry group is exactly of dimension 4. We add the isotropy type in each case, as well as if $\mathfrak{g}$ is an ideal in $\mathfrak{L}$ or not. Also, we include the derived subalgebra of $\mathfrak{L}$, $\mathfrak{L},\mathfrak{L}]$, so that the subsection where each metric appears can be easily identified.
$\mathfrak{g}$ Metric normal forms Isotropy type $\mathfrak{L}$, $\mathfrak{L}$] $\mathfrak{g}$ ideal in $\mathfrak{L}$
@@
$\mathfrak{aff}(\mathbb{R})\oplus\mathbb{R}$ $(e^{1})^{2}+2(e^{2}e^{3})$ hyperbolic $\mathbb{R}^{2}$ yes
@@
Some comments about Table 3 are in order. We start by observing that $\mathfrak{euc}(2)$ does not appear in this table, there are no metrics of 1-dimensional isotropy. The unique metrics on $\mathfrak{psh}$ and on each $\mathfrak{e}(\mu)$ with $\mu>0$ are of constant sectional curvature (zero on $\mathfrak{psh}$ and positive on $\mathfrak{e}(\mu)$). On $\mathfrak{h}(1)$, it is known that all Lorentzian metrics are incomplete, [VS20]. They are all Einstein and apart from the Riemannian (which corresponds to hyperbolic space $\mathbb{H}^{3})$, all have isometry group of dimension 4. It can be seen from Appendix A that all metric normal forms on $\mathfrak{h}(1)$ are present in Tables 2 and 3. A similar situation happens for the Heisenberg group, the list contains all normal forms of metrics. We have a Lorentzian flat space form, and all the other metrics have isometry group of dimension 4. However, all metrics on $\mathfrak{heis}$ are complete, since $\mathfrak{heis}$ is 2-step nilpotent [Gue94].
@@ 2. Preliminaries
Let $(G,g)$ be a 3-dimensional metric Lie group, that is, a Lie group $G$ equipped with a left-invariant (Riemannian or Lorentzian) metric $g$. We will assume that $G$ is connected and simply connected (see [O’N83, Ch. 9, Prop. 20]). Let $L$ denote the full isometry group of $(G,g)$ and $I$ the isotropy subgroup of $L$ at the identity $e_{G}\in G$. Denote also by $\mathfrak{g}$, $\mathfrak{L}$, and $\mathfrak{I}$ the Lie algebras of $G$, $L$ and $I$, respectively. We shall be referring to $\mathfrak{L}$ as the Killing algebra of $(G,g)$ and to its derived subalgebra $\mathfrak{L},\mathfrak{L}]$ as the derived Killing algebra.
@@ 2.1. Infinitesimal Lorentzian isometries
Let $(V,\langle-,-\rangle)$ be a linear Lorentzian space, that is, a vector space $V$ of dimension $n$ equipped with a scalar product of signature (n1,1)(n-1,1).
@@
A linear endomorphism $A:V\longrightarrow V$ is said to be an infinitesimal Lorentzian isometry if $A$ is skew-symmetric with respect to $\langle-,$-$\rangle$, i.e.
@@ 2.1.1. Hyperbolic type
AA has 3 real eigenvalues not all vanishing.
@@ 2.1.2. Parabolic (or nilpotent) type
Then $A$ is nilpotent. It can be readily shown, from the non-degeneracy of $\langle-,$-$\rangle$ that the rank of $A$ cannot be 1. This means that $A$ cannot be 2-nilpotent. Therefore there exists $u\in V$ such that $\{A^{2}(u),A(u),u\}$ is a basis of $V$ and in which the matrix of $A$ is written as $\left(\begin{smallmatrix}0&1&0\\ 0&0&1\\ 0&0&0\end{smallmatrix}\right)$. The exponential of $A$, $\mathrm{exp(A)}$, is unipotent and has a fixed lightlike direction $\ell$ (the span of $A^{2}(u)$) and, therefore, also fixes the orthogonal lightlike plane to $\ell$ (which contains $\ell$).